Hello my name if Jennifer Wheeler and I am confused with my homework. I know some of it but it is just certain problems that I don't get. Can you help me please. Here's a problem
4/4^3=what
I get the problems like for an example: 4^2=1/16 or even 16^0 equals 1.
Pleases help me thankyou

Answer by stanbon(48547)

Here's a problem
4/4^3=what
I get the problems like for an example: 4^2=1/16 or even 16^0 equals 1.
Pleases help me thankyou
4/4^3 means 4/[4*4*4]
A pair of 4's cancel leaving you with 1/[4*4]=1/4^2
By the way: 4^2 does not equal 1/16
4^2 means 4*4 which equals 16
Cheers,
Stan H.